Grade 12Trigonometry

sin120

SEELA YUVARAJU
10 Years agoGrade 12

1 Answer

Vijay Mukati
10 Years ago
Dear Student, sin120 = sin(90+30) = cos30 = sqrt(3)/2. Thanks.